OpenStudy (anonymous):

find the radius of a circle with area 400 pie

OpenStudy (anonymous):

what is the formula for the area of a circle?

OpenStudy (anonymous):

stand aside kiddies. Lien's here to save tha day.

OpenStudy (anonymous):

pie x r^2 ?

OpenStudy (anonymous):

yep. so, \[A=\pi r^2\] you are given \(A=400\pi\) we plug it in \[400\pi=\pi r^2\]solve for "r"

OpenStudy (anonymous):

20?

OpenStudy (anonymous):

bingo

OpenStudy (anonymous):

@studyology Lein's too late

OpenStudy (anonymous):

My food was done. you win this round boy. lml

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

yw